OpenStudy (anonymous):

how do I solve the quadratic equation x squared plus 4 x plus 4 equals 9

OpenStudy (anonymous):

you substract over the 4 then take the 4th root of that to get your x.

OpenStudy (anonymous):

oh wait.

OpenStudy (anonymous):

\[x^2 +4x + 4 = 5\] First way:\[(x+2)^2 = 9\]\[x + 2 = \pm 3\] x = 1 or x = -5 Second way:\[x^2 + 4x - 5 =0\]\[(x+5)(x-1) = 0\] x = 1 or x= -5

OpenStudy (anonymous):

what he said